A car advert says the dealer would give anyone a $500 trade in on his old car. The asking price for the car (before trade in) is $4000, they will offer you a 4% interest rate for a 3year loan. What is the amount you need to finance? Use the pmt function to determine your monthly payment. Copy the work and then use the goal seek tool to determine what the asking price of the car before trade in would have to be to make a monthly payment of $70. Present the payment and loan on a bar chart.


1
Expert's answer
2021-11-18T06:58:24-0500

"A=(4000-500)\cdot(1+\frac{3\cdot 4\%}{100\%})=3920~\$."

"M=\frac{3920}{3\cdot 12}=108.89~\$ ."